OpenStudy (anonymous):

In class i was asked to show my steps of how you would go from cos^2(x)(sec^2x-1) to cot(x). I have no idea how to do this! Can you please show a step by step instructions. (I'm a VERY visual learner!)

OpenStudy (freckles):

do you know Pythagorean identities?

OpenStudy (freckles):

@DanJS you could have left what you had

OpenStudy (danjs):

all good..

OpenStudy (anonymous):

I have the Pythagorean identities on a paper that i can use in class, but to be completely honest, my teacher didn't teach them very well.

OpenStudy (freckles):



OpenStudy (freckles):

you can divide both sides by cos^2(theta) to get the identity in terms of tan and sec

OpenStudy (freckles):

which results in \[1+\tan^2(\theta)=\sec^2(\theta)\]

OpenStudy (freckles):

this is the identity dan had up a sec ago

OpenStudy (danjs):

i was thinking about drawing secant, cosecant lines and such on there, but then no

OpenStudy (freckles):

you can replace sec^2(x) with 1+tan^2(x) in your expression see what this gives you after doing this

OpenStudy (freckles):

if you want you can show us what you have so we know you have done this

OpenStudy (anonymous):

sorry i an super lost, but if you can list all the steps on one post, i can usually figure things out.

OpenStudy (freckles):

Well the first step I gave as: replacing sec^2(x) with 1+tan^2(x) You will see 1-1=0 and be left with cos^2(x)(tan^2(x)) see if you can finish from here

OpenStudy (danjs):

here is the geometry of each trig function, you can see the right triangle for sin^2+cos^2=1^2 inside the circle the other right triangle from tan^2 + 1^2 = sec^2 is the larger bottom right triangle pythagorean theorem

OpenStudy (anonymous):

ok, I'm still having a hard time comprehending all this Trig stuff (it's after midnight where I'm at), but I REALLy appreciate you trying to help. I think that I'll ask my teacher tomorrow.

OpenStudy (danjs):

you need to show \[\cos^2(x)*(\sec^2(x) - 1) = \cot(x)\] ?

OpenStudy (anonymous):

yep!

OpenStudy (danjs):

i dont know if those are equivalent

OpenStudy (jianenriquez):

Anyone else getting \[\sin ^2(x)\] instead of cot(x)

OpenStudy (danjs):

can you check the prob again and see if we have it correct

OpenStudy (anonymous):

oops...totally mixed two questions up!! (sorry everybody!) it should read cos^2(x)(sec^2x-1) is sin^2)(x)

OpenStudy (jianenriquez):

Look at the photo i sent, perhaps you can understand it better

OpenStudy (danjs):

yep, there it is

OpenStudy (jianenriquez):

i forgot to add, \[\tan^2(x) = \frac{ \sin^2(x) }{ \cos^2(x) }\]

OpenStudy (anonymous):

thanks for your work! (sorry a typed everything down wrong!) that's the exact answer i needed!!

OpenStudy (jianenriquez):

heres the edited photo if you need it

OpenStudy (anonymous):

This was perfect!!! Thanks a million!!! :)